OpenStudy (anonymous):

Solve the system by the method of your choice 3x+5y=-18 5x-4y=7 The solution is ? (type an ordered pair)

OpenStudy (anonymous):

4(3x+5y)=-18(4) 5(5-4y)=7(4)

OpenStudy (anonymous):

12x+20y=-72 + 25x-20y=35 ----------------- 37x=-37 x=-1 (sorry tht was supposed to be 5(5X-4Y) and 7(5)

OpenStudy (anonymous):

so if x=-1, plug -1 in for x and you'll get y=-3

OpenStudy (anonymous):

(-1,-3)